During today shift, we permuted
- "rotor R16, box 01" with "rotor R30 box 04"
- "rotor R17, box 6" with "rotor R32 box 08"
with this change, we have R16 on WWN, R17 on WEN, R30 on NWN and R32 on NEN.
We took care to respect (and note) the rotor orientation at all positions. They are as follows:
- with respect to the rotor axis, the LED of the box is on the mirror side for NWN, NNF, NSF, WWN, WEN positions
- the LED is on the opposite site for NEN, NNN, NSN, WSN, WNN
We replaced the photodiode board, LED and box temperature probe for NCal box 04 which had a disconnected wire for the temperature probe.
In the NEB, we replaced the NCal "West" power supply (ID: 3629.6776.04-108784-Ls located at the bottom of the left rack) which had channel 1 and 2 "dead" by a new power supply (ID: 3629.6776.04-106682-Mm). This allowed to restore the NNN power.
we also checked :
- that the twist angles of the exchanged rotors were and still are at 8° with respect to the plates (see picture)
- the height of the "hat" of boxes 1 and 6 is 12 mm (see picture)
- box 4 and 8 do not have any "hat"
Injected signals from permuted NCals were adapated accordingly in NCalmoni and setdefault macros (set_NEB_default_lines.sh and set_WEB_default_lines.sh).
At the end of the operation, all 10 rotors have been successfully restarted at their nominal frequencies with injected lines slightly above 36 Hz.